# Collision Avoidance Verification of Multiagent Systems with Learned Policies

### v1

@@ Abstract
For many multiagent control problems, neural networks (NNs) have enabled promising new capabilities. However, many of these systems lack formal guarantees (e.g., collision avoidance, robustness), which prevents leveraging these advances in safety-critical settings. While there is recent work on formal verification of NN-controlled systems, most existing techniques cannot handle scenarios with more than one agent. To address this research gap, this paper presents a backward reachability-based approach for verifying the collision avoidance properties of Multi-Agent Neural Feedback Loops (MA-NFLs). Given the dynamics models and trained control policies of each agent, the proposed algorithm computes relative backprojection sets by solving a series of Mixed Integer Linear Programs (MILPs) offline for each pair of agents. Our pair-wise approach is parallelizable and thus scales well with increasing number of agents, and we account for state measurement uncertainties, making it well aligned with real-world scenarios. Using those results, the agents can quickly check for collision avoidance online by solving low-dimensional Linear Programs (LPs). We demonstrate the proposed algorithm can verify collision-free properties of a MA-NFL with agents trained to imitate a collision avoidance algorithm (Reciprocal Velocity Obstacles). We further demonstrate the computational scalability of the approach on systems with up to 10 agents.
@@ I Introduction
Verifying the properties of multiagent systems has been an important research area for several decades [1, 2, 3]. Meanwhile, neural network (NN) control policies are becoming a key component of many state-of-the-art multiagent systems, such as for swarming [4] and autonomous driving [5], yet the above verification algorithms cannot deal with these NNs. Obtaining formal guarantees (e.g., for collision avoidance, robustness) for closed-loop systems with NN controllers, i.e., neural feedback loops (NFLs), remains challenging primarily due to the high dimensionality and nonlinearities of NNs. Recent literature aims to provide these formal guarantees, typically by formulating reachability analysis problems and using convex relaxations to obtain tractable verification algorithms [6, 7, 8, 9, 10, 11]. However, existing work focuses primarily on verification of scenarios with a single agent.
Figure 1: Complex interactions between agents present challenges in formal safety verification. This analysis is further complicated when the agents are using NN control policies. To address this issue, this paper provides a backward reachability-based framework for formally verifying collision avoidance of MA-NFLs.
As shown in 1, extending these ideas to multiagent systems raises additional challenges, both in analyzing the complex multiagent interactions and in handling the increase in dimensionality. Ref. [12] is the first approach for formally verifying multiagent NFLs (MA-NFLs). In particular, [12] extends Reach-SDP [8] to compute forward reachable sets over a single representation that contains all agents’ controllers and dynamics. While this approach provides meaningful bounds for several multiagent systems, our experiments suggest that systems with agents trained to avoid collisions are not well-suited for forward reachability-based techniques. For example, in 2, one agent uses the vector field policy from [10], and the other agent is static at the origin (i.e., using a policy that always commands zero velocity). Since the (convex) forward reachable sets (computed by merging both controllers as in [12], then solving mixed integer linear programs (MILPs)) contain both ways the agent could avoid the static agent, the reachable sets intersect with the obstacle, meaning the algorithm is unable to verify collision avoidance. Meanwhile, backward reachability analysis algorithms, such as BReach-LP [10] (which can handle this scenario as a single-agent problem), prove the agents will not collide. However, BReach-LP (and its recent extensions) do not handle more general multiagent systems (e.g., with multiple non-static agents).
@@
Figure 2: (a): Forward reachable sets of agent 1 (blue) intersect with (static) agent 2 (red). We cannot verify safety of system, despite not finding any simulated rollouts (green lines) that enter the obstacle. (b): Result of ReBAR (magenta) is within the obstacle (red), meaning the system is verified safe.
Therefore, this paper aims to extend backward reachability analysis to MA-NFLs. Backward reachability-based algorithms typically compute backprojection sets of the avoid sets [10, 13]. However, this representation is ineffective in the context of collision avoidance for MA-NFLs. Revisiting the scenario in 1, when agents start within some square, the collision set (i.e., when the distance between agents is below a safety threshold) might contain multiple disjoint hyper-rectangles. This representation does not restrict the collisions’ coordinates, and thus computing its backprojection set will result in unbounded answers. In this work, we propose ReBAR and ReBAR-MA, computing relative backprojection sets in a relative coordinate frame using MILPs. This approach enables successful verification of the scenario in 2. In particular, 2(b) shows that the relative backprojection set computed by ReBAR (magenta) lies within the collision set (red), meaning the agents can only collide if they start in a collision. Furthermore, for systems that cannot be verified safe offline, we propose an efficient online safety check using LPs. In summary, the contributions of this work include:
@@
  • demonstrations of ReBAR and ReBAR-MA on relevant multiagent systems, including a 2-agent NFL trained to imitate the Reciprocal Velocity Obstacle (RVO) algorithm [14] and systems with up to 10 agents.
@@ III Approach
III-A Relative Backprojection Sets
@@ IV Experiment
All results are obstained using the optimization solver Gurobi [15] on a PC running Ubuntu 22.04 with a Intel i9-13900K CPU (and an NVIDIA 4090 GPU used during training). ReBAR verifies a pair of agents mimicing RVO [14] offline in 11s and run safety check online in 1.4ms on average. We computed RBPOAs up to 5 steps, and visualize them with the RBPUAs we generated by exhaustive sampling. For systems with up to 10 agents trained to mimic the vector field in [10] with [20,20] hidden neurons, ReBAR-MA verifies each pair of agents in 200s to 320s, and the online check takes less than 1.6ms for each pair of agents.
@@ V Conclusion
This paper presented ReBAR, an algorithm for verifying the collision avoidance safety of Multi-Agent Neural Feedback Loops (MA-NFLs) by computing the relative backprojection set over approximation with state uncertainty offline, and using the result to provide a real-time safety guarantee. Furthermore, we extended ReBAR to provide safety guarantee over an extended horizon, and systems with more than 2 agents in a parallelizable manner (ReBAR-MA). We demonstrate ReBAR and ReBAR-MA by verifying agents trained mimicing Reciprocal Velocity Obstacle (RVO) [14], and system with up to 10 agents. Future work will consider algorithmic extensions to handle general activation functions and dynamics through linear relaxation and verification in decentralized observation space.

### v2

# mybibliography.bib

@@ Abstract
For many multiagent control problems, neural networks (NNs) have enabled promising new capabilities. However, many of these systems lack formal guarantees (e.g., collision avoidance, robustness), which prevents leveraging these advances in safety-critical settings. While there is recent work on formal verification of NN-controlled systems, most existing techniques cannot handle scenarios with more than one agent. To address this research gap, this paper presents a backward reachability-based approach for verifying the collision avoidance properties of Multi-Agent Neural Feedback Loops (MA-NFLs). Given the dynamics models and trained control policies of each agent, the proposed algorithm computes relative backprojection sets by (simultaneously) solving a series of Mixed Integer Linear Programs (MILPs) offline for each pair of agents. We account for state measurement uncertainties, making it well aligned with real-world scenarios. Using those results, the agents can quickly check for collision avoidance online by solving low-dimensional Linear Programs (LPs). We demonstrate the proposed algorithm can verify collision-free properties of a MA-NFL with agents trained to imitate a collision avoidance algorithm (Reciprocal Velocity Obstacles). We further demonstrate the computational scalability of the approach on systems with up to 10 agents.
@@ I Introduction
Verifying the properties of multiagent systems has been an important research area for several decades [pallottino2006probabilistic, doan2014verifying, kouvaros2019formal]. Meanwhile, neural network (NN) control policies are becoming a key component of many state-of-the-art multiagent systems, such as for swarming [tolstaya2020learning] and autonomous driving [palanisamy2020multi], yet the above verification algorithms cannot deal with these NNs. Obtaining formal guarantees (e.g., for collision avoidance, robustness) for closed-loop systems with NN controllers, i.e., neural feedback loops (NFLs), remains challenging primarily due to the high dimensionality and nonlinearities of NNs. Recent literature aims to provide these formal guarantees, typically by formulating reachability analysis problems and using convex relaxations to obtain tractable verification algorithms [dutta2019reachability, huang2019reachnn, hu2020reach, everett2021reachability, rober2022backward, xu2020automatic]. However, existing work focuses primarily on verification of scenarios with a single agent.
Figure 1: Complex interactions between agents present challenges in formal safety verification. This analysis is further complicated when the agents are using NN control policies.
As shown in Fig. 1, extending these ideas to multiagent systems raises additional challenges, both in analyzing the complex multiagent interactions and in handling the increase in dimensionality. Ref. [yan2022strategy, yan2023partially] considers the verification of a two agent system with neuro-symbolic agents. Ref. [gates2023scalable] is the first approach for formally verifying multiagent NFLs (MA-NFLs). In particular, [gates2023scalable] extends Reach-SDP [hu2020reach] to compute forward reachable sets over a single representation that contains all agents’ controllers and dynamics. While this approach provides meaningful bounds for several multiagent systems, our experiments suggest that systems with agents trained to avoid collisions are not well-suited for forward reachability-based techniques, especially when the given controllers are safe. In Fig. 2, one agent uses the vector field policy from [rober2022backward], and the other agent is static at the origin (i.e., using a policy that always commands zero velocity). Since the (convex) forward reachable sets (computed by merging both controllers as in [gates2023scalable], then solving mixed integer linear programs (MILPs)) contain both ways the agent could avoid the static agent, the reachable sets intersect with the obstacle, meaning the algorithm is unable to verify collision avoidance. Meanwhile, backward reachability analysis algorithms, such as BReach-LP [rober2022backward] (which can handle this scenario as a single-agent problem), prove the agents will not collide because although multiple disjoint paths might lead to the target set, these paths must remain within the target set. However, BReach-LP (and its recent extensions) do not handle more general multiagent systems (e.g., with multiple non-static agents).
@@
Figure 2: (a): Forward reachable sets of agent 1 (blue) intersect with agent 2 (red). Unable to verify system safety despite no simulated rollouts (green lines) enters agent 2. (b): Result of ReBAR (magenta) is within agent 2, system is verified safe.
Therefore, this paper aims to extend backward reachability analysis to MA-NFLs. Backward reachability-based algorithms typically compute backprojection sets of the avoid sets [rober2022backward]. However, computing backprojection sets in the global frame is ineffective in the context of collision avoidance for MA-NFLs. Unlike in single agent setting where the avoid set is static, our avoid set is a moving region controlled by another neural network. Computing its backprojection set in the global frame will be thus equivalent to computing the backprojection set of the whole state space. In this work, we propose Relative BAckward Rechability (ReBAR) and ReBAR-MA, computing relative backprojection sets in a relative coordinate frame using MILPs. This approach enables successful verification of the scenario in Fig. 2. In particular, Fig. 2(b) shows that the relative backprojection set computed by ReBAR (magenta) lies within the collision set (red), meaning the agents can only collide if they start in a collision. Furthermore, for systems that cannot be verified safe offline, we propose an efficient online safety check using LPs. In summary, the contributions of this work include:
@@
  • demonstrations of ReBAR and ReBAR-MA on relevant multiagent systems, including a 2-agent NFL trained to imitate the Reciprocal Velocity Obstacle (RVO) algorithm [van2008reciprocal] and systems with up to 10 agents.
@@ III Approach
III-A Relative Backreachable & Backprojection Sets
@@ IV Experiment
All results are obstained using the optimization solver Gurobi [gurobi] on a PC running Ubuntu 22.04 with a Intel i9-13900K CPU. ReBAR verifies a pair of agents mimicing RVO [van2008reciprocal] offline in 11s and run safety check online in 1.4ms on average. We computed RBPOAs up to 5 steps, and visualize them with the RBPUAs we generated by exhaustive sampling. For systems with up to 10 agents trained to mimic the vector field in [rober2022backward] with [20,20] hidden neurons, ReBAR-MA verifies each pair of agents in 200s to 320s, and the online check takes less than 1.6ms for each pair of agents.
@@ V Conclusion
This paper presented ReBAR and ReBAR-MA for verifying the collision avoidance safety of Multi-Agent Neural Feedback Loops (MA-NFLs) by computing the relative backprojection set over approximation with state uncertainty offline, and using the result to provide a real-time safety guarantee. We demonstrate them by verifying agents trained mimicing RVO [van2008reciprocal], and system with up to 10 agents. Future work will consider algorithmic extensions to handle general activation functions and dynamics through linear relaxation and verification in decentralized observation space.
[…]